# Parallel Program Analysis on Path Ranges

### v1

@@ 8.1 Analysis Combinations
Different strategies exist to combine different analyses or analysis tools. For example, one may use an analysis or tool as a component in another approach [32, 33, 34, 35, 36, 37, 38, 39, 40]. Furthermore, one may follow the idea of the strategy design pattern [41] and perform algorithm selection [42] to choose an appropriate analysis from a set of analyses [43, 44, 45, 46, 47, 48, 49]. Other strategies control the execution order of the combination. For instance, the different analyses can be executed sequentially. Sequential portfolio approaches [50, 51] define an order of the analyses and continue with the next analysis if the previous analysis did not succeed, e.g., due to time out or failure. In contrast, testification approaches [52, 53, 7, 54, 55, 56, 57] should make the analysis result more precise. Hence, they typically combine a verification and a validation approach and prioritize or only report confirmed proofs and alarms. To increase the cooperation among sequential combinations even further, some sequential combinations [58, 59, 60, 61, 3, 62, 4, 8, 63, 64, 65] restrict the subsequent analyses to the yet unexplored state space. An extension of sequential combinations is to iteratively run sequential combinations resulting in so called interleaved combinations [66, 67, 68, 6, 9, 58, 69, 70, 71, 72, 73, 74, 75, 76, 77]. Typically, interleaved combinations may exchange some kind of information. An alternative to sequential or interleaved combination is the parallel combination, which we use in our ranged program analysis. The simplest parallel combination is a parallel portfolio [78, 58, 79, 80, 81, 82] that runs different analysis independently in parallel. More refined parallel combinations work together. They may exchange information encountered during analysis [83, 15, 84, 85, 86, 87, 5, 88]. Also, they may distribute the search space among different instances of the same analysis [89, 90, 39, 10, 91, 92, 93, 94, 95, 96, 97, 98]. Our ranged program analyses and instrumentation-based ranged program analysis [24], which encodes the range constraints into the program instead of running a range reduction in parallel to the analysis, go beyond the latter and allows to distribute the search space among different analyses.
@@ 8.2 Search Space Partitioning
To let analyses explore different parts of the search space, the search space needs to be partitioned. Partitioning can be done dynamically based on the already performed search or statically before the exploration. A common dynamic partitioning is to consider the already performed analysis and divide the search space into an already covered and an yet uncovered part. Thereby, the search space may be described based on test goals [58, 59], open proof obligations [60], a previous analysis result [99], program paths [61, 6, 3, 62, 4, 8, 9, 63, 64, 65], or the already explored symbolic execution tree [100, 101, 39, 10, 102, 90]. Static partitioning may for instance divide the analysis into separate subtasks [103, 104, 58] or use a static state partitioning function [105, 106, 107, 108, 109] to divide the states into different sets. Another partitioning strategy is to split the program paths. Conditional static analysis [98] uses the order of executed program branches to characterize the different sets of program paths. Similarly, some ranged symbolic execution approaches [91, 92] [95] characterize the sets of paths via path prefixes. In contrast, one ranged symbolic execution approach [93] defines input constraints to define the sets of paths. Korat [89] defines input ranges while our approach, some ranged symbolic execution approaches [39, 10, 94], and ranged model checking specify path ranges. Ranged symbolic execution creates path ranges from an initial shallow symbolic execution [91, 92, 93] or tests [39, 10, 94]. Our ranged program analysis uses splitting strategies, e.g., based on randomization or loop unrollings. Instrumentation-based ranged program analysis [24] reuses our splitting strategy based on loop unrollings.
@@ 9 Conclusion
Data Availability Statement. All experimental data and our open-source implementation are publicly available in our supplementary artifact [31].
Appendix
Appendix A Additional Figures
A Additional Figures for RQ1
We compare in Sec. 7.2 the increase of the overall execution time for symbolic execution when used as ranged program analysis. In Fig. 13, we provide the remaining figures for predicate abstraction, BMC, and value analysis.
B Additional Figures for RQ2
In Sec. 7.3, we compare the efficiency of ranged program analysis using two different basic analyses. In Fig. 14, we present a more detailed comparison of the overall run time of the basic analyses with the ranged program analysis.

### v2

@@ 8.1 Analysis Combinations
Different strategies exist to combine different analyses or analysis tools. For example, one may use an analysis or tool as a component in another approach [33, 34, 35, 36, 37, 38, 39, 40, 41]. Furthermore, one may follow the idea of the strategy design pattern [42] and perform algorithm selection [43] to choose an appropriate analysis from a set of analyses [44, 45, 46, 47, 48, 49, 50]. Other strategies control the execution order of the combination. For instance, the different analyses can be executed sequentially. Sequential portfolio approaches [51, 52] define an order of the analyses and continue with the next analysis if the previous analysis did not succeed, e.g., due to time out or failure. Testification is the process of looking for evidence to back up the claim that a property is violated or fulfilled. Thereby, testification approaches [53, 54, 7, 55, 56, 57, 58] should make the analysis result more precise. Hence, they typically combine a verification and a validation approach and prioritize or only report confirmed proofs and alarms. To increase the cooperation among sequential combinations even further, some sequential combinations [59, 60, 61, 10, 3, 62, 4, 8, 63, 64, 65] restrict the subsequent analyses to the yet unexplored state space. An extension of sequential combinations is to iteratively run sequential combinations resulting in so called interleaved combinations [66, 67, 68, 6, 9, 59, 69, 70, 71, 72, 73, 74, 75, 76, 77]. Typically, interleaved combinations may exchange some kind of information. An alternative to sequential or interleaved combination is the parallel combination, which we use in our ranged program analysis. The simplest parallel combination is a parallel portfolio [78, 59, 79, 80, 81, 82] that runs different analysis independently in parallel. More refined parallel combinations work together. They may exchange information encountered during analysis [83, 16, 84, 85, 86, 87, 5, 88]. Also, they may distribute the search space among different instances of the same analysis [89, 90, 40, 11, 91, 92, 93, 94, 95, 96, 97, 98]. Our ranged program analyses and instrumentation-based ranged program analysis [26], which encodes the range constraints into the program instead of running a range reduction in parallel to the analysis, go beyond the latter and allow to distribute the search space among different analyses.
@@ 8.2 Search Space Partitioning
To let analyses explore different parts of the search space, the search space needs to be partitioned. Partitioning can be done dynamically based on the already performed search or statically before the exploration. A common dynamic partitioning is to consider the already performed analysis and divide the search space into an already covered and a yet uncovered part. Thereby, the search space may be described based on test goals [59, 60], open proof obligations [61], a previous analysis result [99], program paths [10, 6, 3, 62, 4, 8, 9, 63, 64, 65], or the already explored symbolic execution tree [100, 101, 40, 11, 102, 90]. Static partitioning may for instance divide the analysis into separate subtasks [103, 104, 59] or use a static state partitioning function [105, 106, 107, 108, 109] to divide the states into different sets. Another partitioning strategy is to split the program paths. Conditional static analysis [98] uses the order of executed program branches to characterize the different sets of program paths. Similarly, some ranged symbolic execution approaches [91, 92] [95] characterize the sets of paths via path prefixes. In contrast, one ranged symbolic execution approach [93] defines input constraints to define the sets of paths. Korat [89] defines input ranges while our approach, some ranged symbolic execution approaches [40, 11, 94], and ranged model checking specify path ranges. In contrast to path ranges, the set of paths described by an input range is not guaranteed to be ordered. Thus, the technique for range reduction presented may not work. Ranged symbolic execution creates path ranges from an initial shallow symbolic execution [91, 92, 93] or tests [40, 11, 94]. Our ranged program analysis uses splitting strategies, e.g., based on randomization or loop unrollings. Instrumentation-based ranged program analysis [26] reuses our splitting strategy based on loop unrollings.
@@ 9 Conclusion
Our experiments suggest that combining a technique whose strength is finding property violations with another one good at computing proofs within ranged program analysis and employing work stealing can form a successful combination that can increase the overall effectiveness. Validating whether novel combinations using program instrumentation for off-the-shelf verifiers also benefit from using work stealing is one of the challenges that we aim to address next. For instance, using complete tools like Esbmc [112], Mopsa [113], or UltimateAutomizer [114] instead of individual analyses in combination seems promising.
Beneath work stealing the generation of the program ranges can influence the performance of ranged program analysis. Currently, we employ Lb3, a splitter that generates the ranges using a fixed heuristic. Finding a more sophisticated technique that may also take the given task into account is a promising line of work, e.g. by using constant values extracted from the program or estimating the number of loop iterations for computing more balanced ranges. Recently, the idea of dynamic, demand-driven splitting is realized in the tool Bubaak-SpLit [115]. Instead of computing the ranges up-front, Bubaak-SpLit generates ranges in case configurations of Bubaak fail to solve tasks within certain limits. This idea opens another promising line of work for ranged program analysis.
Lastly, the information computed by one ranged analysis may also be helpful for the other analysis running in parallel. Hence, we aim to establish information exchange among ranged analyses, to enable cooperation within ranged program analysis. Beneath that, we aim to evaluate if generating more than one counterexample within ranged program analysis is beneficial.
Data Availability Statement. All experimental data and our open-source implementation are publicly available in our supplementary artifact [25]. Therein, we also included a virtual machine for reproducing the experiments.